OpenStudy (anonymous):

What is the solution set of the following equation? -8x + x + 15 = -7x + 12

OpenStudy (anonymous):

can you simplify the equation? it could be easier for you to find the set when it is simplified

OpenStudy (anonymous):

what do you do to simplify?

OpenStudy (anonymous):

combined like terms, but with solution sets you don't combine across the equal sign. I'll show you. combining the -8x and x will give us -7x. Now if we think about this, is there any number in the universe that could make this equation true? -7x + 15 = -7x + 12

OpenStudy (anonymous):

umm 3?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

no there is no number that will make it true. Let's take away the -7 because there is one on each side of the equal sign and we are left with: x + 15 = x + 12 There really is no number we could substitute for x that would make this true right?